a sample of 10.5 g of nitrogen reacts with 20.2 g of hydrogen to produce ammonia


1
Expert's answer
2021-02-15T03:24:08-0500

N2 +3H2 "\to" 2NH3

no. Of moles of N2 = 10.5g/28gmol- = 0.375 mole

no. Of moles of H2 = 20.2g/2gmol- = 10.1 mole

Based on mole ratio, N2 is limiting in this situation because there is more than enough H2 but not enough N2.

moles NH3 can be produced = (0.375 moles N2 × 2 moles NH3 )/1 mole N2 = 0.75 moles

grams of NH3 = 0.75 × 17 = 12.75 g NH3 produced